The equation x^2/(2-k)+y^2/(k-5)+1=0 represents an ellipse , if
(a)k<5
(b)k<2
(c)2(d)k<2 or k>5

Grade:

1 Answers

masu mounika
11 years ago

answer

x^2/(2-k)+y^2/(k-5)=-1

x^2/(k-2)+y^2/(5-k)=1

where "a"is the semi major axis a^2=k-2

k=a^2+2   (k>2)

"b" is semi minor axis b^2=5-k

k=5-b^2     (k<5)

so answer is (a)